OpenStudy (anonymous):

Convert the numeral to a numeral in the base indicated. 16,629 to base 13

hartnn (hartnn):

u know a method to corvert a decimal number to other base number?

OpenStudy (anonymous):

oops.. sorry.. keep dividing the number by 13 and note down remainders.. and taking remainders from down to up.. right hexadecimal equivalents of them :P

OpenStudy (anonymous):

16629 / 13 = 1279 1279 * 13 = 16627 16629 - 16627 = 2 16629 in base 13 is 7752 ?

OpenStudy (anonymous):

damn i am sorry !! :(

OpenStudy (anonymous):

? :D wait I'm not right :(

hartnn (hartnn):

@vikrantg4 whats hexadecimal equivalent ?? u copied that from somewhere ?

hartnn (hartnn):

your first step is correct @Esmy , u get the remainder=2

hartnn (hartnn):

2nd step also correct, u get remainder = 5

hartnn (hartnn):

3rd step correct, remainder = 7

OpenStudy (anonymous):

I'm I correct over all :'D

hartnn (hartnn):

yup.

OpenStudy (anonymous):

YAYAYAY :'D thanks ^-^

hartnn (hartnn):

welcome ^_^

OpenStudy (amistre64):

another method would be to write down the exponential values of a base 13 system, and work it from there 13^4 13^3 13^2 13^1 13^0 28561 2197 169 13 1

OpenStudy (anonymous):

Ohh :o okay that seems cool. Ill try that one in my next one :)

OpenStudy (amistre64):

:) just to refresh my memory .... 16,629 is divisible by 13^3: 7 times , 1250 left over 1250 is divisible 13^2: 7 times , 67 left over 67 is divisible 13^1: 5 times , 2 left over \[16,629_{10}=7752_{13} \]